# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

### v1

@@ 5.4 Performance Summary
Table 5 presents a comparative analysis of the proposed ADFSA FS method against conventional techniques including PCA and RFE across four widely used classifiers: SVM, RF, KNN, and LogReg. The evaluation is based on three key metrics: cross-validation accuracy, number of selected features, and average inference time over five runs. The ADFSA method consistently achieves the highest accuracy across all classifiers, reaching up to 99.65% for RF, while significantly reducing the feature set to only seven features. This demonstrates the method’s capability to retain highly discriminative features while discarding redundant or irrelevant ones. Additionally, ADFSA achieves the lowest inference times in all cases, indicating its suitability for real-time human activity recognition applications. The combined improvements in accuracy, feature compactness, and computational efficiency clearly highlight the superiority of the proposed ensemble fitness function over traditional FS methods.
@@
Figures 10 and 11 illustrate a comprehensive comparative evaluation of different FS methods applied to SVM and RF classifiers, respectively. The results clearly highlight the superior performance of the proposed ADFSA approach in all metrics considered. For the SVM classifier, ADFSA achieves the highest accuracy of 99.58%, outperforming traditional methods such as PCA and RFE, which yield 97.79% and 97.73%, respectively. Notably, ADFSA also uses only 7 features, significantly fewer than PCA (50) and the baseline (128), demonstrating its efficiency in dimensionality reduction. Additionally, it reports the lowest inference time of 0.196 ms, indicating suitability for real-time scenarios.
@@
Similarly, in the case of the RF classifier (Figure 11), the proposed method achieves the best accuracy of 99.65%, compared to 97.45% with PCA and 97.51% with RFE. It also maintains a reduced feature set of 7 while lowering the inference time to 8.714 ms, which is significantly faster than the baseline method (19.189 ms). These results affirm that ADFSA not only preserves or improves classification performance but also reduces computational overhead, making it highly effective for Human Activity Recognition tasks where both accuracy and real-time efficiency are crucial.
@@
Figures 12 and 13 further validate the effectiveness of the proposed ADFSA FS method when applied to KNN and LogReg classifiers. For the KNN classifier (Figure 12), ADFSA achieves the highest accuracy of 99.58%, outperforming PCA (97.54%) and RFE (97.44%). More importantly, it reduces the number of features to just 7, compared to 50 for PCA/RFE and 128 in the baseline. Inference time also shows a significant improvement, dropping to 6.958 ms, whereas PCA and RFE report 8.033 ms and 17.676 ms, respectively. This performance reinforces the scalability of ADFSA even in non-parametric models like KNN.
@@
Similarly, Figure 13 shows that in the case of LogReg, ADFSA achieves 99.51% accuracy, significantly higher than the baseline (97.39%), PCA (97.58%), and RFE (97.57%). It also demonstrates the most efficient inference time at 0.173 ms, with only 7 selected features. In contrast, PCA and RFE require 50 features each and yield higher inference times of 0.966 ms and 0.202 ms, respectively. These results highlight the proposed method’s strong generalization ability across different classifier types while consistently reducing computational cost and improving classification performance.
Figure 14 shows the confusion matrices of the selected classifiers; all four classifiers exhibit excellent discrimination capability across the 11 activity classes. In each case, the confusion matrix shows strong diagonal dominance, indicating that the majority of the test samples were correctly classified. Notably, there are no significant misclassifications across classes, and only a few isolated off‐diagonal entries can be observed. This consistent behaviour across SVM, Random Forest, KNN, and Logistic Regression demonstrates the effectiveness of the proposed feature extraction and selection strategy in capturing highly discriminative representations for HAR.
@@
(c) KNN
(d) Logistic Regression
Figure 14: Confusion Matrices
@@ 5.5 Analysis of Memory Usage and Training Time
Table 7 provides a comparative view of memory usage and training time for different classifiers under various feature selection strategies. It can be observed that the proposed ADFSA method consistently achieves lower memory consumption and faster training compared to traditional approaches such as PCA and RFE. For instance, SVM and Logistic Regression show a remarkable drop in both resource usage and training duration with ADFSA, while KNN benefits significantly in terms of memory reduction. Although Random Forest maintains a similar memory footprint across methods, ADFSA still contributes to a noticeable decrease in training time. Overall, the table highlights that the proposed approach offers a more efficient balance between memory usage and computational cost across diverse classifiers.
@@
Table 8 highlights the impact of various FS techniques on memory consumption and training time across four classifiers. The baseline (No FS) represents the unaltered scenario, against which percentage changes are compared. The proposed ADFSA method consistently demonstrates the most significant reductions in both memory usage and training time. For instance, in the case of SVM and KNN, ADFSA achieves memory reductions of 84.75% and 90.96%, respectively, along with training time decreases exceeding 66%. Similarly, for LogReg, ADFSA lowers memory usage by 83.53% and nearly eliminates training time (–98.06%), emphasizing its efficiency. In contrast, traditional methods such as PCA and RFE show mixed outcomes: although PCA significantly reduces memory and training time for most classifiers, it slightly increases memory usage in RF. RFE, while effective in reducing memory, results in increased training time in SVM and offers limited efficiency gains in other classifiers. These results reinforce the capability of ADFSA to enhance resource efficiency, making it particularly suitable for real-time and resource-constrained machine learning environments.
@@
Figures 15 to 18 present a comparative assessment of memory consumption and training time across four classifiers, SVM, RF, KNN, and LogReg when applied with different FS techniques.
@@
The effectiveness of ADFSA over memory usage and training time can be clearly seen from Figures 15 and Figure 16 of both the classifiers SVM and RF, respectively. The reduction in both memory usage and training time for the SVM classifier can be seen in Figure 15, about an 84.75% drop in memory and a 79.11% decrease in training time achieved by the proposed ADFSA scheme as compared to PCA and RFE. It is therefore observed that the unusability of RFE for the SVM classifier in this regard. In contrast, all the methods have nearly occupied constant memory across all techniques, but still proposed ADFSA observing a decrease of 62.91% in training time, as compared to only 5.07% and 29.08% achieved by PCA and RFE, respectively. Overall, it is clearly observing the robustness and efficacy of the proposed ADFSA in terms of FS in real-time and computationally constrained environments.
@@ 5.6 Other Performance Matrices
Table 9 presents the comparative performance of four classifiers, SVM, RF, KNN, and LogReg across different FS strategies, namely no FS, PCA, RFE, and the proposed ADFSA method. For all classifiers, the proposed ADFSA approach consistently achieved the highest F1-score, precision, and recall values, demonstrating a significant performance improvement over the baseline and conventional FS methods. Specifically, SVM reached an F1-score of 99.58%, RF attained 99.65%, KNN achieved 99.58%, and Logistic Regression obtained 99.51% with ADFSA, surpassing other configurations by notable margins. In contrast, the No FS, PCA, and RFE settings showed relatively close but lower results, typically in the 97.2%–97.8% range for all metrics. These findings indicate that ADFSA not only enhances classification accuracy but also maintains a balanced trade-off between precision and recall, highlighting its effectiveness in improving overall model performance.
@@
The statistical analysis for F1-score is summarized in Table 10, which reports the F1-scores with 95% confidence intervals and corresponding p-values for each classifier under different feature selection settings. The proposed ADFSA method consistently achieves the highest F1-scores across all classifiers, with values exceeding 99.5% and narrow confidence intervals, indicating superior accuracy and stability. In contrast, baseline configurations without feature selection, as well as PCA- and RFE-based approaches, yield F1-scores in the range of 97.2–97.8%, with differences statistically significant at p << 0.001. These results demonstrate the robustness and generalizability of ADFSA in enhancing classification performance compared to conventional feature selection techniques.
Overall, the analysis confirms that ADFSA consistently enhances computational efficiency across diverse classifiers. While traditional techniques like PCA and RFE contribute to some extent, they fall short in achieving the balance between memory economy and training speed that ADFSA provides. This reinforces the suitability of ADFSA as a versatile FS strategy for scalable machine learning pipelines.
Figures 19-22 present a comparative analysis of the performance metrics, F1-score, Precision, and Recall for four classifiers: SVM, RF, KNN, and LogReg, respectively. Each figure includes grouped bar charts that visualize the impact of four FS methods: No FS, PCA (n=50), RFE, and the proposed ADFSA technique.
@@
The results clearly indicate that across all classifiers, the ADFSA method consistently achieves the highest scores in all three evaluation metrics. For example, Figure 19 shows that SVM with ADFSA reaches an F1-score of 99.58%, while Figure 20 highlights a similar peak of 99.65% for the RF classifier. The consistent upward trend across Figure 21 and Figure 22 further confirms the superiority of the ADFSA method when applied to KNN and LogReg. These improvements across different learning algorithms demonstrate the robustness and generalizability of the proposed FS strategy.
@@ 5.7 Ablation Study
Figure 23 and Table 11 presents an ablation study conducted using the SVM classifier to evaluate the individual and combined effects of the uniqueness-based reward and complexity penalty in the proposed FS strategy. Three configurations were tested: (i) No Uniqueness, No Complexity Penalty, (ii) No Complexity Penalty, and (iii) Full (both components enabled).
@@
Figure 24 and 25 illustrates the impact of different fitness function configurations on the optimizer’s weight update behavior and FS performance. In the first setting, Fig. 24(a), where neither uniqueness nor complexity penalties are applied, the optimizer enters a stagnation phase after just 6 generations, resulting in the selection of 52 features. This early stagnation indicates poor exploration capability and a tendency to converge prematurely. In the second configuration, Fig. 24(b), which includes only the complexity penalty, the optimizer shows a modest improvement by stagnating after 12 generations and selecting 64 features, suggesting some control over model complexity but still lacking diversity in the solution space. In contrast, the third configuration, Fig. 25 uses the complete fitness function with both uniqueness and complexity components. This setup allows the optimizer to utilize the full range of generations without premature stagnation and leads to a significantly reduced and optimal feature subset of only 7 features. These results demonstrate the importance of balancing exploration and exploitation through well-designed fitness objectives in FS tasks.
@@ 5.8 Comparison of Proposed Framework with the State Of The Art
Recent transformer-based and hybrid models have further improved spatial–temporal learning; for example, ViT-ReT [21], 3DCNN [18], and VGG-BiGRU [20] achieved 92.4%, 85.20%, and 93.38% accuracy, respectively. More advanced approaches, such as the 2024 bidirectional LSTM [22] and ConvNeXt-TCN [19], reported 99.2% and 97.73%. By comparison, the proposed architecture yields a higher accuracy of 99.58% on the same 80/20 split by effectively integrating spatial–temporal feature learning with an optimized feature selection strategy, demonstrating strong generalization and suitability for real-time HAR applications.

### v2

@@ 5.3 Performance Summary
Table 5 presents a comparative analysis of the proposed ADFSA FS technique against conventional techniques including PCA and RFE across four widely used classifiers: SVM, RF, KNN, and LogReg. The evaluation is based on three key metrics: cross-validation accuracy, number of selected features, and average inference time over five runs. The ADFSA method consistently achieves the highest accuracy across all classifiers, reaching up to 99.65% for RF, while significantly reducing the feature set to only seven features. This demonstrates the method’s capability to retain highly discriminative features while discarding redundant or irrelevant ones. Additionally, ADFSA achieves the lowest inference times in all cases, indicating its suitability for real-time human activity recognition applications. The combined improvements in accuracy, feature compactness, and computational efficiency clearly highlight the superiority of the proposed ensemble fitness function over traditional FS techniques.
@@
Figures 10 and 11 illustrate a comprehensive comparative evaluation of different FS methods applied to SVM and RF classifiers, respectively. The results clearly highlight the superior performance of the proposed ADFSA technique in all metrics considered. For the SVM classifier, ADFSA achieves the highest accuracy of 99.58%, outperforming traditional methods such as PCA and RFE, which yield 97.79% and 97.73%, respectively. Notably, ADFSA also uses only 7 features, significantly fewer than PCA (50) and the baseline (128), demonstrating its efficiency in dimensionality reduction. Additionally, it reports the lowest inference time of 0.196 ms, indicating suitability for real-time scenarios.
@@
Similarly, in the case of the RF classifier (Figure 11), the proposed technique achieves the best accuracy of 99.65%, compared to 97.45% with PCA and 97.51% with RFE. It also maintains a reduced feature set of 7 while lowering the inference time to 8.714 ms, which is significantly faster than the baseline method (19.189 ms). These results affirm that ADFSA not only preserves or improves classification performance but also reduces computational overhead, making it highly effective for Human Activity Recognition tasks where both accuracy and real-time efficiency are crucial.
@@
Figures 12 and 13 further validate the effectiveness of the proposed ADFSA FS technique when applied to KNN and LogReg classifiers. For the KNN classifier (Figure 12), ADFSA achieves the highest accuracy of 99.58%, outperforming PCA (97.54%) and RFE (97.44%). More importantly, it reduces the number of features to just 7, compared to 50 for PCA/RFE and 128 in the baseline. Inference time also shows a significant improvement, dropping to 6.958 ms, whereas PCA and RFE report 8.033 ms and 17.676 ms, respectively. This performance reinforces the scalability of ADFSA even in non-parametric models like KNN.
@@
Similarly, Figure 13 shows that in the case of LogReg, ADFSA achieves 99.51% accuracy, significantly higher than the baseline (97.39%), PCA (97.58%), and RFE (97.57%). It also demonstrates the most efficient inference time at 0.173 ms, with only 7 selected features. In contrast, PCA and RFE require 50 features each and yield higher inference times of 0.966 ms and 0.202 ms, respectively. These results highlight the strong generalization of the proposed technique in different types of classifiers while consistently reducing computational cost and improving classification performance.
Figure 14(a) and Figure 14(b) present the confusion matrices for SVM and Random Forest classifiers, respectively. Both classifiers exhibit strong discrimination capability across the 11 activity classes. In each case, the confusion matrix demonstrates clear diagonal dominance, indicating that the majority of the test samples were correctly classified. Only a few isolated off-diagonal entries are present, suggesting minimal misclassifications between activity classes.
@@
Figure 14: Confusion Matrices of SVM and Random Forest
Figure 15(a) and Figure 15(b) illustrate the confusion matrices for the KNN and Logistic Regression classifiers. Similar to SVM and Random Forest, both classifiers achieve consistent recognition performance with strong diagonal dominance across the activity classes. No major misclassifications are observed, and only a limited number of off-diagonal entries occur. These results further validate the effectiveness of the proposed feature extraction and selection technique in capturing highly discriminative representations for human activity recognition.
(a) KNN
(b) Logistic Regression
Figure 15: Confusion Matrices of KNN and Logistic Regression
For clarity, the correspondence between the class indices (0–10) and their respective activity labels is given in Table 7. This consistent behaviour across SVM, Random Forest, KNN, and Logistic Regression highlights the effectiveness of the proposed feature extraction and selection technique in capturing highly discriminative representations for HAR.
Table 7: Class Activity Labels.
Index Activity Index Activity
0 Basketball 6 Swing
1 Biking 7 Tennis Swing
2 Diving 8 Trampoline Jumping
3 Golf Swing 9 Volleyball Spiking
4 Horse Riding 10 Walking
5 Soccer Juggling
@@ 5.4 Analysis of Memory Usage and Training Time
Table 8 provides a comparative view of memory usage and training time for different classifiers under various FS strategies. It can be observed that the proposed ADFSA technique consistently achieves lower memory consumption and faster training compared to traditional approaches such as PCA and RFE. For instance, SVM and Logistic Regression show a remarkable drop in both resource usage and training duration with ADFSA, while KNN benefits significantly in terms of memory reduction. Although Random Forest maintains a similar memory footprint across methods, ADFSA still contributes to a noticeable decrease in training time. Overall, the table highlights that the proposed technique offers a more efficient balance between memory usage and computational cost across diverse classifiers.
@@
Table 9 highlights the impact of various FS techniques on memory consumption and training time across four classifiers. The baseline (No FS) represents the unaltered scenario, against which percentage changes are compared. The proposed ADFSA technique consistently demonstrates the most significant reductions in both memory usage and training time. For instance, in the case of SVM and KNN, ADFSA achieves memory reductions of 84.75% and 90.96%, respectively, along with training time decreases exceeding 66%. Similarly, for LogReg, ADFSA lowers memory usage by 83.53% and nearly eliminates training time (–98.06%), emphasizing its efficiency. In contrast, traditional methods such as PCA and RFE show mixed outcomes: although PCA significantly reduces memory and training time for most classifiers, it slightly increases memory usage in RF. RFE, while effective in reducing memory, results in increased training time in SVM and offers limited efficiency gains in other classifiers. These results reinforce the capability of ADFSA to enhance resource efficiency, making it particularly suitable for real-time and resource-constrained machine learning environments.
@@
The effectiveness of ADFSA on memory usage and training time can be clearly seen from Figures 16 and Figure 17 of both the SVM and RF classifiers, respectively. The reduction in both memory usage and training time for the SVM classifier can be seen in Figure 16, about an 84.75% drop in memory and a 79.11% decrease in training time achieved by the proposed ADFSA technique as compared to PCA and RFE. It is therefore observed that the unusability of RFE for the SVM classifier in this regard. In contrast, all the methods have nearly occupied constant memory across all techniques, but still proposed ADFSA observing a decrease of 62.91% in training time, as compared to only 5.07% and 29.08% achieved by PCA and RFE, respectively. Overall, it is clearly observed that the robustness and efficacy of the proposed ADFSA in terms of FS in real-time and computationally constrained environments.
@@ 5.5 Other Performance Matrices
Table 10 presents the comparative performance of four classifiers, SVM, RF, KNN, and LogReg across different FS strategies, namely no FS, PCA, RFE, and the proposed ADFSA technique. For all classifiers, the proposed ADFSA technique consistently achieved the highest F1-score, precision, and recall values, demonstrating a significant performance improvement over the baseline and conventional FS methods. Specifically, SVM reached an F1-score of 99.58%, RF attained 99.65%, KNN achieved 99.58%, and Logistic Regression obtained 99.51% with ADFSA, surpassing other configurations by notable margins. In contrast, the No FS, PCA, and RFE settings showed relatively close but lower results, typically in the 97.2%–97.8% range for all metrics. These findings indicate that ADFSA not only enhances classification accuracy but also maintains a balanced trade-off between precision and recall, highlighting its effectiveness in improving overall model performance.
@@
The statistical analysis for F1-score is summarized in Table 11, which reports the F1-scores with 95% confidence intervals and corresponding p-values for each classifier under different FS settings. The proposed ADFSA technique consistently achieves the highest F1-scores across all classifiers, with values exceeding 99.5% and narrow confidence intervals, indicating superior accuracy and stability. In contrast, baseline configurations without FS, as well as PCA- and RFE-based approaches, yield F1-scores in the range of 97.2–97.8%, with differences statistically significant at p << 0.001. These results demonstrate the robustness and generalizability of ADFSA in enhancing classification performance compared to conventional FS techniques.
@@
Furthermore, Figures 20-23 present a comparative analysis of the performance metrics, including F1-score, Precision, and Recall, for four classifiers: SVM, RF, KNN, and LogReg, respectively. Each figure includes grouped bar charts that visualize the impact of four FS methods: No FS, PCA, RFE, and the proposed ADFSA technique.
The results clearly indicate that across all classifiers, the ADFSA method consistently achieves the highest scores in all three evaluation metrics. For example, Figure 20 shows that SVM with ADFSA reaches an F1-score of 99.58%, while Figure 21 highlights a similar peak of 99.65% for the RF classifier. The consistent upward trend across Figure 22 and Figure 23 further confirms the superiority of the ADFSA method when applied to KNN and LogReg. These improvements across different learning algorithms demonstrate the robustness and generalizability of the proposed FS technique.
@@
Overall, the analysis confirms that ADFSA consistently enhances computational efficiency across diverse classifiers. Although traditional techniques like PCA and RFE contribute to some extent, they fail to achieve the balance between memory economy and training speed that ADFSA provides. This reinforces the suitability of ADFSA as a versatile FS strategy for scalable machine learning pipelines.
@@ 5.6 Ablation Study
Figure 24 and Table 12 presents an ablation study conducted using the SVM classifier to evaluate the individual and combined effects of the uniqueness-based reward and complexity penalty in the proposed FS technique. Three configurations were tested: (i) No Uniqueness, No Complexity Penalty, (ii) No Complexity Penalty, and (iii) Full (both components enabled).
@@
Figure 25 and 26 illustrates the impact of different fitness function configurations on the optimizer’s weight update behavior and FS performance. In the first setting, Fig. 25(a), where neither uniqueness nor complexity penalties are applied, the optimizer enters a stagnation phase after just 6 generations, resulting in the selection of 52 features. This early stagnation indicates poor exploration capability and a tendency to converge prematurely. In the second configuration, Fig. 25(b), which includes only the complexity penalty, the optimizer shows a modest improvement by stagnating after 12 generations and selecting 64 features, suggesting some control over model complexity but still lacking diversity in the solution space. In contrast, the third configuration, Fig. 26 uses the complete fitness function with both uniqueness and complexity components. This setup allows the optimizer to utilize the full range of generations without premature stagnation and leads to a significantly reduced and optimal feature subset of only 7 features. These results demonstrate the importance of balancing exploration and exploitation through well-designed fitness objectives in FS tasks.
@@ 5.7 Comparison of Proposed Framework with the State Of The Art
Recent transformer-based and hybrid models have further improved spatial–temporal learning; for example, ViT-ReT [21], 3DCNN [18], and VGG-BiGRU [20] achieved 92.4%, 85.20%, and 93.38% accuracy, respectively. More advanced approaches, such as the 2024 bidirectional LSTM [22] and ConvNeXt-TCN [19], reported 99.2% and 97.73%. By comparison, the proposed framework yields a higher accuracy of 99.58% on the same 80/20 split by effectively integrating spatial–temporal feature learning with an optimized FS strategy, demonstrating strong generalization and suitability for real-time HAR applications.